# Accelerating Vision Diffusion Transformers with Skip Branches

### v2

@@ Abstract
Diffusion Transformers (DiT), an emerging image and video generation model architecture, has demonstrated great potential because of its high generation quality and scalability. Despite the impressive performance, its practical deployment is constrained by computational complexity and redundancy in the sequential denoising process. While feature caching across timesteps has proven effective in accelerating diffusion models, its application to DiT is limited by fundamental architectural differences from U-Net-based approaches. Through empirical analysis of DiT feature dynamics, we identify that significant feature variation between DiT blocks presents a key challenge for feature reusability. To address this, we convert standard DiT into Skip-DiT  with skip branches to enhance feature smoothness. Further, we introduce Skip-Cache  which utilizes the skip branches to cache DiT features across timesteps at the inference time. We validated the effectiveness of our proposal on different DiT backbones for video and image generation, showcasing skip branches to help preserve generation quality and achieve higher speedup. Experimental results indicate that Skip-DiT achieves a 1.5×1.5$\times$1.5 × speedup almost for free and a $\times2$ speedup with only a minor reduction in quantitative metrics. Code is available at https://github.com/OpenSparseLLMs/Skip-DiT.git.
Figure 1: Results and generation speed comparison of Skip-DiT  cached with Skip-Cache  and its original version. Utilizing the skipping branch, Skip-DiT can significantly boost inference speed while preserving original quality. Latency is measured on one A100. We use Latte as the backbone model for video generation, and Hunyuan-DiT for image generation.
@@ 1 Introduction
Diffusion models [9, 3, 24, 46] have emerged as the de-facto solution for visual generation, owing to their high fidelity outputs and ability to incorporate various conditioning signals, particularly natural language. Classical diffusion models, which adopt U-Net [27] as their denoising backbone, have dominated image and video generation applications. More recently, Diffusion Transformers (DiT) [4, 23] have introduced an alternative architecture that replaces traditional sequential convolutional networks with Vision Transformers, offering enhanced scalability potential. While initially designed for image generation, DiT has demonstrated remarkable effectiveness when extended to video generation tasks [19, 16, 25]. However, despite these advances, significant challenges remain in scaling diffusion models efficiently, particularly for applications involving large numbers of input tokens such as video generation. This scaling challenge is especially pronounced in DiT architectures, where the attention mechanism’s computational complexity grows quadratically with input size. The magnitude of this challenge is illustrated by Sora [20], a state-of-the-art video generation model that remains unavailable to the public.
Numerous approaches have been proposed to enhance the efficiency of diffusion models, including reduced sampling techniques [33], distillation methods [41, 29], and quantization strategies [5]. Caching mechanisms, which reuse noise latent across timesteps, have emerged as a particularly promising direction as they do not require extensive model retraining [18, 15, 38, 47, 6]. However, many existing caching approaches [18, 38] are specifically tailored to U-Net architectures, leveraging their unique structural properties—particularly the skip connections between paired downsampling and upsampling blocks that enable high-level feature reuse while selectively updating low-level features. While some recent studies [47, 6] have attempted to adapt caching mechanisms for DiT acceleration, they have not achieved the same level of efficiency gains and performance preservation as their U-Net counterparts.
(a) Denoising step = 10
(b) Denoising step = 250
Figure 2: Feature smoothness analysis of DiT with class-to-video generation on Taichi using DDPM. Normalized disturbances (with strength coefficients $\alpha$ and $\beta$) are added to the model with and without skip connections. We compare the similarity between the original and perturbed features. The feature difference surface of Latte, with and without skip connections, is visualized in steps 10 and 250 of DDPM.
To understand the key challenges of feature caching in DiT, we analyze the feature dynamics during the denoising process. Drawing inspiration from loss landscape visualization techniques [13, 11], we examine feature changes using the early and late timesteps in denoising as a case study. For effective caching, features should exhibit minimal variation between timesteps, allowing us to reuse features from previous steps and bypass computation in subsequent Transformer blocks. We term this property ”feature smoothness”, which manifests as flatness in the landscape visualization. However, as illustrated in Figure 2, we observe that vanilla DiT (w/o skip branches) exhibits high feature variance across timesteps, contrary to the desired characteristics for effective caching. Thus, we ask:
(Q) What are the new insights can we gain from feature smoothness in DiT architectures, and how can these insights inform the design of more effective inference caching mechanisms?
Drawing from the findings in [13], where residual connections are shown capable of mitigating sharp loss landscapes, in this study, ❶ we first conduct preliminary experiments adding skip branches to pre-trained DiT model from shallow to deep blocks, named Skip-DiT . which achieves significantly improved feature smoothness with minimal continuous pre-training, as demonstrated in Figure 2 (w/ skip branches). ❷ We then leverage these skip branches during inference to implement an efficient caching mechanism, Skip-Cache , where only the first DiT block’s output needs to be computed for subsequent timesteps while deep block outputs are cached and reused. ❸ To evaluate our proposal, we conduct extensive experiments across multiple DiT backbones, covering image and video generation, class-conditioned and text-conditioned generation. We demonstrate that Skip-DiT consistently outperforms both dense baselines and existing caching mechanisms in both qualitative and quantitative evaluations. Our contributions are three-fold:
  • We identify feature smoothness as a critical factor limiting the effectiveness of cross-timestep feature caching in DiT, which helps better understand caching efficiency.
  • We build Skip-DiT , a skip-branch augmented DiT architecture that enhances feature smoothness, and Skip-Cache that efficiently leverages skip branches for feature caching across timesteps. Skip-Cache facilitates accelerated inference via caching while maintaining visual generation performance.
  • Extensive empirical evaluations demonstrate that Skip-Cache achieves substantial acceleration: up to 1.5×1.5$\times$1.5 × speedup is achieved almost for free and a $\times2$ speedup with only a minor reduction in quantitative metrics.
@@ 2 Related Works
The diffusion model has become the dominating architecture for image and video generation, whose main idea is iterative generate high-fidelity images or video frames from noise [26]. Early diffusion models mainly employ U-Net as their denoising backbone [24, 3]. However, U-Net architectures struggle to model long-range dependencies due to the local nature of convolutions. Researchers proposing diffusion transformer model (DiT) for image generation [4, 2, 23]. Recent years have witnessed a significant growth in studies of video DiT. Proprietary DiT such as Sora [20] and Movie-Gen [25] show impressive generation quality, also evidenced by open-sourced implementation [48, 12]. Latte decomposes the spatial and temporal dimensions into four efficient variants for handling video tokens, allowing effective modeling of the substantial number of tokens extracted from videos in the latent space [19]. CogvideoX adds a 3D VAE combined with an expert transformer using adaptive LayerNorm, which enables the generation of longer, high-resolution videos [40]. However, as the number of tokens grows exponentially with video length and spatial resolution, the computational complexity of DiT especially the self-attention mechanism remains a significant bottleneck for video generation.
@@
Since the diffusion model involves iterative denoising, caching features across time-steps, model layers, and modules has been found an effective way to save inference computation costs. For U-Net Diffusion, DeepCache [18] and FRDiff [32] exploit temporal redundancy by reusing features across adjacent denoising steps. While other works take a more structured approach by analyzing and caching specific architectural components–Faster Diffusion [15] specifically targets encoder feature reuse while enabling parallel decoder computation, and Block Caching [38] introduces automated caching schedules for different network blocks based on their temporal stability patterns. Recently, cache-based acceleration has also been applied to DiT. PAB [47] introduces a pyramid broadcasting strategy for attention outputs. $\Delta$-DiT [6] proposes adaptive caching of different DiT blocks based on their roles in a generation–rear blocks during early sampling for details and front blocks during later stages for outlines. T-Gate [45] identifies a natural two-stage inference process, enabling the caching and reuse of text-oriented semantic features after the initial semantics-planning stage. While these caching techniques have shown promise, they are primarily limited to inference-time optimization, and there remains significant potential for improving their acceleration factors.
@@ 3 Methodology
Figure 3: Illustration of Skip-DiT  and Skip-Cache for DiT visual generation caching. (a) The vanilla DiT block for image and video generation. (b) Skip-DiT modifies the vanilla DiT model using skip branches to connect shallow and deep DiT blocks. (c) Given a Skip-DiT  with $L$ layers, during inference, at the $t-1$ step, the first layer output $x^{\prime t-1}_{0}$ and cached $L-1$ layer output $x^{\prime t}_{L-1}$ are forwarded through the skip branches to the final DiT block to generate the denoising output, without executing DiT blocks 2 to $L-1$.
@@ 3.1 Preliminaries
In our study, we consider two types of DiT models processing different visual information: image DiT and video DiT, presented in Figure 3 (a). ❶ Image DiT model follows Vision Transformer, each block contains self-attention, cross-attention, and feed-forward networks (FFN), where the cross-attention module incorporates text and timestep conditions. ❷ Video DiT adopts dual-subblock architecture: the spatial Transformer subblock processes tokens within the same time frame, while the temporal subblock manages cross-frame relationships. These blocks alternate in sequence, with cross-attention integrating conditional inputs. A complete video DiT block pairs one spatial and one temporal component in an interleaved pattern, following [19, 47].
3.2 Visualizing the Feature Smoothness of DiT
Section 1 introduces the concept of feature smoothness, and provides an intuitive explanation. In this section, we will detail the feature smoothness visualization and analysis.
To use this approach, we denote the original module parameters of the base model as $\theta^{*}$ in the graph and choose two random direction vectors,$\delta$ and $\eta$, each direction vectors share the same dimension as $\theta$. Firstly, these directions are normalized according to the original parameters it correspond to. We take $\delta$ and $\eta$ to disturb the model with strength coefficients $\alpha$ and $\beta$. After updating, we get a new model with parameter $\theta^{{}^{\prime}}$:
@@ 3.2 Visualizing the Feature Smoothness of DiT
$\theta^{{}^{\prime}}=\theta^{*}+\alpha\delta+\beta\eta$ (3)
We denote the predicted noise after $k$ denoising steps of model before and after adding disturbs as $\bm{x}_{\theta^{*}}^{k}$ and $\bm{x}_{\theta^{{}^{\prime}}}^{k}$. We then define the feature difference with function;
$L(\theta^{*})=\frac{\bm{x}_{\theta^{*}}^{k}\cdot\bm{x}_{\theta^{{}^{\prime}}}^{k}}{\|\bm{x}_{\theta^{*}}^{k}\|\|\bm{x}_{\theta^{{}^{\prime}}}^{k}\|}$ (4)
we then plot the feature difference surface feature according to the 2-D function:
$f(\alpha,\beta)=L(\theta^{*}+\alpha\delta+\beta\eta)$ (5)
This approach was employed in [8] and [14], where $L(\theta^{*})$ represents the loss of a model with parameters $\theta^{*}$, used to analyze trajectories of various minimization methods and model structures. Similarly, [11, 13] utilized this approach to demonstrate that different optimization algorithms converge to distinct local minima in the 2D projected space.
3.3 Skip-DiT: Improving Feature Smoothness
We visualize vanilla DiT feature smoothness in Figure 2 (w/o Skip branches) which is trained and generated on Taichi dataset. We can observe drastic feature changes at the two DDPM steps, indicating they are not ideal states to cache features, thus shortening the space for caching. According to the insights from [13] and Diffusion caching study utilizing U-Net residual connection feature [18], we next investigate whether DiT feature smoothness can be improved by minimal modification to its structure to insert residual property, i.e. skip branches.

### v4

# Towards Stabilized and Efficient Diffusion Transformers through Long-Skip-Connections with Spectral Constraints

@@ Abstract
Diffusion Transformers (DiT) have emerged as a powerful architecture for image and video generation, offering superior quality and scalability. However, their practical application suffers from inherent dynamic feature instability, leading to error amplification during cached inference. Through systematic analysis, we identify the absence of long-range feature preservation mechanisms as the root cause of unstable feature propagation and perturbation sensitivity. To this end, we propose Skip-DiT, an image and video generative DiT variant enhanced with Long-Skip-Connections (LSCs) - the key efficiency component in U-Nets. Theoretical spectral norm and visualization analysis demonstrate how LSCs stabilize feature dynamics. Skip-DiT architecture and its stabilized dynamic feature enable an efficient statical caching mechanism that reuses deep features across timesteps while updating shallow components. Extensive experiments across the image and video generation tasks demonstrate that Skip-DiT achieves: (1) $\times4$ training acceleration and faster convergence, (2) 1.52×1.5-2$\times$1.5 - 2 × inference acceleration with negligible quality loss and high fidelity to the original output , outperforming existing DiT caching methods across various quantitative metrics. Our findings establish Long-Skip-Connections as critical architectural components for stable and efficient diffusion transformers. Codes are provided in the URL [^1] https://github.com/OpenSparseLLMs/Skip-DiT.
Figure 1: (a) Feature similarities between standard and cache-accelerated outputs in vanilla DiT (caching with FORA [30] and Faster-Diff [13]) and our proposed Skip-DiT. Skip-DiT presents consistently higher feature similarity, demonstrating superior stability after caching. (b) Illustration of Skip-DiT that modifies vanilla DiT models using long-skip-connection to connect shallow and deep DiT blocks. Dashed arrows indicate paths where computation can be skipped in cached inference. (c) Comparison of video generation quality (PNSR) and inference speedup of different DiT caching methods. Skip-DiT maintains higher generation quality even at greater speedup factors.
@@ 1 Introduction
Diffusion models [8, 3, 24, 51] have emerged as the de-facto solution for visual generation, owing to their high fidelity outputs and ability to incorporate various conditioning signals, particularly natural language. Classical diffusion models typically adopt U-Net [27] as their denoising backbone. Recently, Diffusion Transformers (DiT) [4, 23] introduce an alternative architecture to replace traditional convolutional networks with Vision Transformers, offering enhanced scalability potential. While initially designed for image generation, DiT has demonstrated remarkable effectiveness when extended to video generation tasks [18, 14, 25]. Despite these advances, DiT still faces challenges in practical deployment due to its slow convergence rates [52, 44, 20] and substantial inference time requirements, posing significant constraints on training efficiency and real-time applications.
Numerous approaches have been proposed to improve the efficiency of diffusion models, including reduced sampling techniques [33], distillation methods [45, 29], and quantization strategies [5]. Among these approaches, caching has emerged as one of the most effective strategies for enhancing inference efficiency, owing to its low computational cost while maintaining high fidelity and similarity to the original models through efficient feature reuse [30, 49, 6, 11, 53, 16]. However, existing caching methods in DiT still face challenges in visual generation. DiT architecture exhibits significant sample-wise variation in caching function responses, requiring adaptive per-sample hyperparameter tuning and introducing computational overhead [11, 16]. Besides, current caching methods still yield suboptimal generative results with visible artifacts when cached with larger intervals [30, 6, 53, 11], suggesting error amplification in DiT’s transformer blocks. Besides, the training efficiency gap further compounds these challenges. DiT is shown to converge more slowly in training [52, 44], and the reasons remain unexplained from architectural perspectives.
DiT efficiency is hindered by the above challenges, but U-Net-based diffusion models can leverage their Long-Skip-Connections (LSCs) to extend cache intervals without complex timestep scheduling or severe performance degradation [17, 13]. The LSCs not only make caching stable and robust during inference, but they have also been proven as the key factor for stabilizing training and accelerating convergence for U-Net-based diffusion models [9]. These observations lead to our core research question:
(Q) What are the architectural root causes of feature instability and inefficiency in DiT, and to what extent can Long-Skip-Connections mitigate these issues while preserving generation quality?
In this study, we systematically analyze the feature distribution in DiT: ❶ We conduct a preliminary experiment that incorporates Long-Skip-Connections (LSCs) with DiT, and visualize the dynamic features of DiT with and without LSCs by adding perturbations in parameters and model inputs, the feature similarity landscape and plots validate the instability of vanilla DiT (w/o LSCs), which we formally characterize as Dynamic Feature Instability, indicating an uncontrolled spectral norm that affects both statistical stability, model robustness, and convergence rate. ❷ We then theoretically prove the superior stability and robustness of DiT with LSCs over vanilla DiT via spectral analysis.
To this end, we propose the Long-Skip-Connected-DiT, namely Skip-DiT, a DiT architecture that employs spectral constraints—limiting the maximum singular value of weight matrices to ensure stable gradient flow and reduced sensitivity to perturbations. By incorporating these constraints through LSCs, our approach provides theoretical feature stability guarantees and enhanced efficiency in both training and inference, validated through quantitative experiments and visual analysis. The LSCs also enable output caching of stabilized features from deep blocks, requiring the computation of only shallow blocks during inference at statically chosen caching timesteps.
To evaluate our proposed Skip-DiT, we conduct extensive experiments across 3 DiT backbones, assess training performance on 6 models, and inference efficiency on 7 visual generation tasks, spanning both class and text conditional image/video generation. Skip-DiT  achieves superior training efficiency while consistently outperforming both standard baselines and prior caching methods in qualitative and quantitative evaluations. To summarize, we claim the following contribution of this work:
  • Through theoretical analysis and comprehensive visualization, we identify Dynamic Feature Instability as a fundamental root cause that impairs DiT training efficiency and hinders effective inference acceleration.
  • We introduce Skip-DiT, an enhanced DiT architecture incorporating Long-Skip-Connections between shallow and deep transformer blocks. Our approach provides theoretical guarantees for feature stability through spectral norm analysis and enables efficient feature caching at strategically selected timesteps.
  • We conduct extensive evaluations across different DiT base models, and various image and video generation benchmarks. Results of substantial training and inference speedup and nearly undamaged generation quality consistently endorse the effectiveness of Skip-DiT. For example, compared to vanilla DiT, Skip-DiT  achieves up to $\times4$ training acceleration with even better generation quality. When compared to state-of-the-art DiT caching techniques, Skip-DiT  delivers $1.5$-2×2$\times$2 × additional speedup while maintaining superior output fidelity.
@@ 2 Related Works
The diffusion model has become the dominating architecture for image and video generation, whose main idea is iterative generate high-fidelity images or video frames from noise [26]. Early diffusion models mainly employ U-Net as their denoising backbone [24, 3]. However, U-Net architectures struggle to model long-range dependencies due to the local nature of convolutions. Researchers proposing diffusion transformer model (DiT) for image generation [4, 2, 23]. Recent years have witnessed a significant growth in studies of video DiT. Proprietary DiT such as Sora [20] and Movie-Gen [25] show impressive generation quality, also evidenced by open-sourced implementation [54, 12]. Latte decomposes the spatial and temporal dimensions into four efficient variants for handling video tokens, allowing effective modeling of the substantial number of tokens extracted from videos in the latent space [18]. CogvideoX adds a 3D VAE combined with an expert transformer using adaptive LayerNorm, which enables the generation of longer, high-resolution videos [43].
@@
Since the diffusion model involves iterative denoising, caching features across time-steps, model layers, and modules has been found an effective way to save inference computation costs. For U-Net Diffusion, DeepCache [17] and FRDiff [32] exploit temporal redundancy by reusing features across adjacent denoising steps. While other works take a more structured approach by analyzing and caching specific architectural components–Faster Diffusion [13] specifically targets encoder feature reuse while enabling parallel decoder computation, and Block Caching [41] introduces automated caching schedules for different network blocks based on their temporal stability patterns. Recently, cache-based acceleration has also been applied to DiT. PAB [53] introduces a pyramid broadcasting strategy for attention outputs. $\Delta$-DiT [6] proposes adaptive caching of different DiT blocks based on their roles in a generation–rear blocks during early sampling for details and front blocks during later stages for outlines. T-Gate [50] identifies a natural two-stage inference process, enabling the caching and reuse of text-oriented semantic features after the initial semantics-planning stage. To deal with the sample-wise variation in DiT, adaptive caching methods like AdaCache [11] and TeaCache [16] are proposed to predict the feature distribution according to the input.
Long Skip Connections in DiT
While Long Skip Connections (LSCs) have been incorporated into Diffusion Transformers (DiTs) by previous works like HunyuanDiT [14] and U-DiT [37], the underlying benefits of these connections have been largely unexamined. This paragraph clarifies the distinct advantages of our LSC implementation. Hunyuan DiT[14] is the first text-to-image model with LSCs, but a systematic exploration of their functional role has been absent from the literature. U-DiT [37] reshapes the Transformer into a U-Net structure [27] for efficiency, using token and feature downsampling for complexity reduction. In contrast, the primary motivation for Skip-DiT  is to improve feature stability and training convergence by using LSCs to optimize the spectral norm. This core design difference leads to a novel statistic-cached inference capability not present in other models. Furthermore, our feature fusion mechanism differs fundamentally. While U-DiT concatenates features, thereby doubling their dimension, Skip-DiT  employs a dimension-preserving fusion strategy that mixes recomputed and cached features from deep-to-shallow layers, enhancing robustness. Crucially, Skip-DiT  offers superior model compatibility. Its architecture, built upon native DiT blocks, allows for the direct adaptation of pretrained models. This is a significant advantage over U-DiT, which requires a full retraining process. For instance, we demonstrate this flexibility by successfully converting a pretrained DiT into our text-to-video model through continued training.
(a) Feature similarity landscape on class-to-image task.
(b) Feature similarity landscape on text-to-video task.
Figure 2: Feature stability comparison between Skip-DiT (left) and vanilla DiT (right) on class-to-image (DiT-XL) and text-to-video (Latte) generation. We inject perturbations $\delta$ (magnitudes $\alpha$) and $\eta$ (magnitudes $\beta$), normalized with specific coefficients ($\epsilon=1e^{-3}$ for DiT-XL and $2e^{-2}$ for Latte), then measure the similarity between the standard and perturbed features.
@@ 3.1 Preliminaries
(a) Class-to-image Task.
(b) Text-to-video Task.
Figure 3: Comparison between standard and cache-accelerated outputs in vanilla DiT versus Skip-DiT, with Latte (text-to-video generation) and DiT-XL (class-to-image generation) serving as base architectures. Both mean and standard deviation across samples are shown. Vanilla DiT exhibits much higher sample variance.
@@
In our study, we consider two types of DiT models processing different visual information: image DiT and video DiT. ❶ Image DiT model follows Vision Transformer, each block contains self-attention, cross-attention, and feed-forward networks (FFN), where the cross-attention module incorporates text and timestep conditions. ❷ Video DiT adopts dual-subblock architecture: the spatial Transformer subblock processes tokens within the same time frame, while the temporal subblock manages cross-frame relationships. These blocks alternate in sequence, with cross-attention integrating conditional inputs. A complete video DiT block pairs one spatial and one temporal component in an interleaved pattern, following [18, 53].
3.2 Dynamic Feature Instability in DiT
DiT models are known to suffer from low training efficiency [52] and unstable feature distributions during inference [11, 16], posing significant challenges for acceleration and their application in interleaved reasoning [36, 35]. This contrasts sharply with U-Net [27] architectures where Long-Skip-Connections (LSCs) enhance training stability and efficiency [9], and enable simple but reliable acceleration through feature reuse [17, 13]. To explore this architectural discrepancy, we integrate LSCs into DiT-XL [22] as in Figure 2 (b) and train from scratch as the preliminary experiment. Training settings is the same as in Section 4.1
Visualizing Feature Stability of DiT
Let $\theta\in\mathbb{R}^{d}$ denote the original parameters of model $f(x;\theta)$. Given random perturbation vectors $\delta,\eta\in\mathbb{R}^{d}$, $\|\delta\|=\|\eta\|=\epsilon\|\theta\|$, with $10<\epsilon<$ 1 which is normalization coefficient. we construct perturbed parameters: $\theta^{\prime}=\theta+\alpha\delta+\beta\eta$, where $\alpha,\beta\in\mathbb{R}$ control perturbation magnitudes. The feature stability landscape for 3D visualization is defined as:
@@ 3.2 Dynamic Feature Instability in DiT
$L(\alpha,\beta)=\mathrm{CosineSimilarity}\left(f(x;\theta),f(x;\theta^{\prime})\right)$ (3)
To visualize the feature stability across timesteps, we perturb the inputs with cached values, following the caching strategy of FORA[30] and FasterDiff[44], and we take cosine similarity to measure the distance with and without caching. Visualized results are illustrated in Figure 2(a) and 3(a).
Dynamic Feature Instability
Our preliminary experiments and visualizations reveal that vanilla DiT models exhibit significant sensitivity to perturbations in both parameters and input features, with feature similarity across timesteps demonstrating considerable variability across different examples. These observations raise critical concerns regarding the stability of DiT models, particularly for training efficiency and inference acceleration. We formally define the instability of DiT as Dynamic Feature Instability (DFI). We preliminary attribute these issues to the uncontrolled spectral norm $\sigma_{\max}$ of DiT’s Jacobians $J^{\text{vanilla}}$, which exponentially amplifies perturbations and destabilizes feature propagation. This aligns with findings in [46], where high spectral norms degrade model robustness and generalization capacity by amplifying input sensitivities, and mirrors the instability patterns observed in visual generative GAN training [19] without spectral normalization.
Training Convergence Comparison
To validate the theoretical convergence rate in real implementations, we extend preliminary experiments recording the relationship between training steps and performance and FID[21] on ImageNet[7]. Both DiT models with/without LSCs are trained from scratch under identical settings. Results are revealed in Figure 4. The accelerated convergence of DiT w/ LSCs observed in practical training demonstrates the training efficiency and model capability. This mirrors our theoretical analysis in Theorem 1, confirming that the spectral norm is effectively controlled with LSCs during training.
Figure 4: Training efficiency comparison between Skip-DiT  and DiT-XL. Skip-DiT achieves superior FID-50K score on ImageNet at 1.6M training steps (vs. DiT-XL’s 7M steps) and converges faster at 2.9M steps. Both models are evaluated on FID-50K under identical settings (no classifier-free guidance, cfg=1).
3.3 Integerating Long-Skip-Connection into DiT